Janet kim is now 13 years old, her father is now 31
write an equation stating that jeong is twice as old as janet, then solve it to find out when jeong is twice as old as janet

Answer by ptaylor(2198)

First off, I'm assuming that Jeong is Janet's father.
Now Janet is 13 and Jeong is 31
Let's say that in x years from now Jeong will be twice as old as Janet
In x years from now, Janet will be 13+x years old
Jeong will be 31+x years old which has to be twice as old as Janet (2(13+x))
So our equation to solve is:
31+x=2(13+x) simplifying, we get:
31+x=26+2x collecting like terms, we will have:
31-26=2x-x or
x=5 years
ck
2(13+5)=31+5
2(18)=36
36=36
Hope this helps. Happy holidays.-----ptaylor